# Attention Is Not What You Need: Revisiting Multi-Instance Learning for Whole Slide Image Classification

### v1

@@ 4.3 Results on Real-World Dataset
Table  1(a) shows the results of our method and other comparative methods on Camelyon16. For patch-level classification, FocusMIL and CausalMIL achieve significantly better results than the attention mechanism-based MIL. During the training process, we observed that the attention-based MIL approach achieved a high patch AUC at the beginning of training, but then gradually declined as the slide AUC increased, as shown in Figure 4. This suggests that the model learns to identify salient instances initially, but later relies on bag context information to fit to hard-to-classify slides. It is worth mentioning that DSMIL achieves better results compared to other attention-based methods. This improvement is attributed to DSMIL’s first stream using a max-pooling based instance-level classifier (Li, Li, and Eliceiri 2021); however, its second-stage attention network allows it to learn bag context information, which reduces its ability to recognize tumour concepts.
When the feature extractor is switched to CTransPatch, the patch-level classification performance of the attention-based MIL methods significantly improves but is still much worse than FocusMIL. This is because ResNet-18 is pre-trained on ImageNet and retains context patterns (Geirhos et al. 2018) in its features. In contrast, CTransPatch is pre-trained on WSI slides with self-supervised contrast learning, and its features help to reduce the dependency of attention-based models on bag context information.
For slide-level classification, FocusMIL significantly outperforms the former best-performing method, DTFD-MIL, in accuracy by $3.43\%$ with Resnet18 and $0.94\%$ with CTransPath, while slightly underperforming in AUC by $0.23\%$ and $0.08\%$, respectively.
Table 2: Performance Comparison on the TCGA Dataset.
Model Slide AUC Slide Acc
ABMIL 0.8900 (CI: 0.8654, 0.9146) 0.8200 (CI: 0.8036, 0.8364)
DSMIL 0.8909 (CI: 0.8849, 0.8970) 0.8295 (CI: 0.8154, 0.8436)
TransMIL 0.9501 (CI: 0.9456, 0.9547) 0.8886 (CI: 0.8778, 0.8993)
DTFD-MIL 0.9340 (CI: 0.9328, 0.9352) 0.8695 (CI: 0.8663, 0.8727)
IBMIL 0.9401 (CI: 0.9390, 0.9411) 0.8743 (CI: 0.8675, 0.8811)
CausalMIL 0.9298 (CI: 0.9256, 0.9339) 0.8791 (CI: 0.8654, 0.8927)
Ours 0.9360 (CI: 0.9309, 0.9411) 0.8819 (CI: 0.8742, 0.8896)
Table 2 presents the slide-level classification results on the TCGA-NSCLC dataset. The performance of FocusMIL is slightly below that of TransMIL but is comparable to DTFD-MIL, IBMIL, and CausalMIL. It is worth mentioning that the initial intention of proposing FocusMIL was to develop a reliable WSI classification method that performs well in tumour region localization and is unaffected by spurious correlations. These bag-based MIL models may exploit spurious correlations between bag context and labels to achieve better slide-level classification on some datasets, but the models will no longer be reliable when this correlation changes. Additionally, TransMIL does not provide a method for predicting patch labels.
@@
CausalMIL 1.000 0.839 (0.823, 0.855) 0.373 (0.333, 0.412)
Ours 1.000 0.852 (0.835, 0.868) 0.402 (0.379, 0.425)
4.4 Results on Standard-MIL test
As shown in Table 3, all methods fit well in the training set. All of the tested attention-based methods have a test slide AUC less than 0.5, failing to respect the standard MIL assumption. CausalMIL and FocusMIL are little affected by the poison. Benefiting from the max-pooling based instance classifier, DSMIL performs well in patch-level classification, further demonstrating that max-pooling is crucial for learning positive concepts.
Model Para. Time C16 ACC TCGA ACC
ABMIL 657k 3.35s 0.8152 0.8200
DSMIL 856k 4.92s 0.7984 0.8295
TransMIL 2.66M 9.56s 0.8127 0.8886
DTFD-MIL 987k 6.02s 0.8032 0.8695
CausalMIL 302k 3.28s 0.8328 0.8791
Ours 167k 3.23s 0.8375 0.8819
4.5 Computational Cost Analysis
As shown in Table 4, FocusMIL has significantly fewer parameters compared to attention-based methods. Due to the use of mini-batch gradient descent with a batch size of 3, both CausalMIL and FocusMIL achieve superior training speeds. However, CausalMIL requires additional computations for reconstructing loss and the KL divergence loss of the conditional prior distribution, which results in slightly longer training times compared to FocusMIL. Although TransMIL shows better performance in the slide-level classification task on the TCGA-NLCSC dataset, its model size and training time are approximately 16 and 3 times greater than ours, respectively. Moreover, TransMIL could not provide patch-level predictions.
Table 5: Ablation study on Camelyon16 Dataset. Dropout was set at 0.4 for the second model.
@@ 4.5 Computational Cost Analysis
FocusMIL 0.860 $\pm$ 0.020 0.411 $\pm$ 0.035
@@
4.6 Ablation Study
As shown in Table 5, the baseline model mi-Net is very unstable. In our experiments, we found that mi-Net may achieve high training slide-level AUC but extremely low patch-level AUCPR during training and fails completely during testing. This severe overfitting issue disappeared when the dropout was set above 0.4. However, this “training failure” reappeared when training mi-Net using mini-batch gradient descent, since establishing correct classification boundaries for salient and hard instances within a batch of multiple slides is more difficult. FocusMIL does not incorporate dropout; instead, forward amortized variational inference helps to avoid “training failure” and achieves a slide AUC that is 1.4% higher than mi-Net with a dropout of 0.4. Further significant improvements of 4.6% in slide AUC are obtained by training multiple bags together in a mini-batch. Since our method does not rely on bag-inherited information $\bm{Z}_{B}$ for classification, a higher slide AUC actually indicates better classification boundaries for hard instances, which is also confirmed by visualization results. For ablation study on TCGA-NSCLC dataset and experiments on the impact of batch size, please refer to the supplementary material.
4.7 Visualization
@@
For fairness of comparison, the models for visualization are not deliberately selected, and both FocusMIL with a training batch size of 3 and DTFD-MIL models achieved a test slide AUC of approximately 0.87. As shown in Figure 5, our model accurately identifies both large and small tumour regions without any false positive predictions. FocusMIL with a training batch size of 1 misses some small tumour regions. For DTFD-MIL, it misses the tumour region in the first slide and some tumour patches in the second slide, and there are many false positive predictions. Notably, in the fourth slide, DTFD-MIL assigns high scores to many patches irrelevant to the tumour. After zooming in, the tissues in these regions are obviously blurrier, which is due to the slide preparation process. This confirms that DTFD-MIL can rely on bag-inherited information for predictions. We provide additional visualization results in the supplementary material.
5 Conclusion
This paper introduces FocusMIL, a method that respects the standard multiple instance learning assumption and incorporates variational inference to ensure the model focuses on tumour morphology. We further utilize mini-batch gradient descent to achieve better classification boundaries for identifying hard instances. Quantitative experiments demonstrate that our method significantly outperforms attention-based methods in patch-level classification and achieves competitive results at the slide-level compared to the current state-of-the-art. The visualization results demonstrate the superiority of our method in tumour region localization.
Broader Impact
The theoretical analysis presented in this paper indicates that attention-based methods cannot guarantee predictions based on the morphology of tumour cells (causal factor). Our quantitative experiments and visualizations also demonstrate that these methods fail to provide reliable tumour region predictions. To make matters worse, results from the Camelyon16 standard MIL test dataset show that attention-based methods can fail when there exists a strong spurious correlation between the bag context and the labels. Therefore, we emphasise that attention-based MIL methods should be used cautiously in WSI classification.
FocusMIL is a significantly more reliable alternative that focuses on causal factors for prediction. We recommend practitioners start with FocusMIL as a solid foundation to avoid excessive risk in deployment. Also, researchers can design more sophisticated WSI classification methods that satisfy the MIL hypothesis based on our FocusMIL.
Appendix A Additional Quantitative Experiments
A.1 Ablation Study on Hyperparameter
Ablation Study on Batch Size
Figure 6 shows the impact of different batch sizes on the model’s slide and patch level classification performance. Batch size greater than 1 can significantly improve the model’s slide-level classification performance. The improvement in patch F1-score is not pronounced. This is due to the fact that salient instances account for a rather high proportion of the overall positive instances, and the model’s ability to recognize hard instances will not be obviously reflected in an increase of the patch F1-score metric. A batch size that is too large can reduce the performance of the model, and we recommend choosing a batch size between 2 and 5.
Figure 6: Results of ablation experiments on batch size.
Ablation Study on Dimentions of Latent Representation
As shown in Figure 7, our method is not sensitive to the parameter of latent dimension.
Figure 7: Results of ablation experiments on dimentions of latent representation
A.2 Ablation Study on TCGA-NSCLC Dataset
Since the positive instance ratio within the positive bags is high in this dataset, mi-Net did not encounter the “training failure” issues that occurred on the Camelyon16 dataset. With the introduction of forward amortized variational inference, FocusMIL improves 2.2% over mi-Net on the slide accuracy metric, and also improves 1.4% over mi-Net with dropout. Mini-batch gradient descent further improves FocusMIL’s performance by 1% on the slide accuracy metric. This improvement is smaller compared to the Camelyon16 dataset, possibly because the morphological differences in tumor cells are not as pronounced in the TCGA dataset as in the Camelyon16 dataset. Since the TCGA-NSCLC dataset does not have patch-level labels, this hypothesis is speculative. It is worth noting that mean-pooling outperforms max-pooling for aggregating slide-level predictions. However, we recommend that mean-pooling should be used with caution because mean-pooling does not respect the standard MIL assumption.
Table 6: Ablation study on TCGA-NSCLC dataset. Dropout was set at 0.4 for the second and third models. For the last two models, mean-pooling is used to aggregate patch-level predictions into slide-level predictions during test.
Model Mini-batch Slide AUC Slide ACC
mi-Net $0.904\pm 0.008$ $0.838\pm 0.023$
mi-Net+dropout $0.913\pm 0.009$ $0.846\pm 0.015$
mi-Net+dropout $0.910\pm 0.009$ $0.850\pm 0.016$
FocusMIL $0.916\pm 0.011$ $0.860\pm 0.016$
FocusMIL $0.919\pm 0.007$ $0.871\pm 0.005$
FocusMIL (mean) $0.936\pm 0.004$ $0.882\pm 0.006$
FocusMIL (mean) $0.934\pm 0.005$ $0.879\pm 0.007$
Appendix B Additional Visualization
As shown in figure 8, DTFD-MIL generates false positive predictions for some patches of normal slides. After zooming in on these patches, we find that these patches that are given high positive probabilities are very blurry compared to the surrounding patches, likely due to bias introduced during the whole slide image production process. This confirms that attention-based methods may make predictions based on spurious correlations, leading to unreliable tumor localization. FocusMIL does not have any false-positive predictions for these normal slides
@@ Appendix B Additional Visualization
Figure 9 demonstrates the visualization results of FocusMIL and DTFD-MIL for five tumor slides. FocusMIL with a training batch size of 3 perfectly predicts all tumour regions. When the training batch size was changed to 1, FocusMIL misses some small tumour regions. This result confirms that training with mini-batch gradient descent does help to establish better classification boundaries for salient and hard instances. For DTFD-MIL, its predictions for both large and small tumour areas are suboptimal due to the fact that it can overfit on salient instances.

### v2

@@ 6.2 Results on Real-World Dataset
Results on Camelyon16 As shown in Table 1, for patch-level classification, max-pooling-based methods such as mi-Net, CausalMIL, and FocusMIL achieve significantly better results than the attention mechanism-based MIL methods. For slide-level classification, mi-Net and CausalMIL achieve performance comparable to ABMIL. Our proposed FocusMIL achieves better results than the more sophisticated attention-based methods in terms of both AUC and ACC. For tumour region localization (Table 2), FocusMIL outperforms DTFD-MIL substantially. It is worth mentioning that after using the mini-batch gradient optimization strategy, FocusMIL achieves better small-area tumour (hard instance) prediction compared to training with SGD. We provide extensive visualization results in the Appendix.
When using features extracted by CTransPath, all models show a noticeable performance boost. In this case, the max-pooling-based methods still achieve significantly better results at the patch-level. For slide-level, three max-pooling-based methods achieve slightly leading results. Notably, DSMIL achieves better results compared to other attention-based methods. This improvement is attributed to DSMIL’s first stream using a max-pooling-based instance classifier.
Results on Camelyon17 Since the training and testing data come from different hospitals, the distribution of the bag context information $\bm{Z}_{B}$ may vary. In this OOD scenario, as shown in Table 3, FocusMIL achieves significantly better results. However, mi-Net and CausalMIL overfit the dataset by memorizing instance features within the bag, leading to training failure. After using features extracted by the advanced foundation model, the three max-pooling-based methods still exhibit slightly better performance.
6.3 Results on Standard-MIL test
As shown in Table 4, all methods fit well in the training set. All of the tested attention-based methods have a test slide AUC less than 0.5, failing to respect the standard MIL assumption. The three max-pooling-based methods are little affected by the poison. Benefiting from the max-pooling-based instance classifier, DSMIL performs well in patch-level classification, further demonstrating that max-pooling is crucial for learning positive concepts.
For the ablation study on FocusMIL, and the effect of batchsize and KL coefficients on the model performance, and computational cost analysis, please refer to Appendix.
@@ 6.3 Results on Standard-MIL test
mi-Net 1.000 0.805 (0.772, 0.838) 0.384 (0.317, 0.453)
CausalMIL 1.000 0.808 (0.785, 0.830) 0.366 (0.325, 0.407)
FocusMIL 1.000 0.868 (0.853, 0.883) 0.384 (0.366, 0.401)
7 Conclusion
In this paper, we analyze the fact that attention-based MIL methods are inherently prone to spurious correlation in WSI classification. Furthermore, we empirically investigate the limitations of existing max-pooling-based MIL methods and introduce FocusMIL, a simple yet effective max-pooling approach. Our comprehensive experiments confirm that max-pooling-based approaches can achieve superior results to the attention-based MIL approaches and exhibit significantly more robust performance in out-of-distribution scenarios. We therefore hope our work will motivate a shift from attention-based methods and a renewed focus on max-pooling-based methods to achieve more robust, interpretable, and trustworthy clinical applications.
Impact Statement
Our paper demonstrates that attention-based MIL methods are often not reliable in computer-aided diagnosis based on histopathology images. Although advanced foundation models can significantly boost their performance, attention-based aggregation approaches remain untrustworthy for high-stakes cancer diagnosis.
In real-world scenarios, the training data and deployment data often originate from different hospitals, resembling our experimental setup on the Camelyon17 dataset. In such conditions, max-pooling-based methods exhibit a significant advantage by focusing on the morphological features of cancer cells while being much less susceptible to spurious correlations. We believe that our position—refocusing on max-pooling-based MIL methods that adhere to the standard multi-instance assumption—will drive computer-aided WSI diagnostic models towards more reliable outcomes.
Furthermore, our results stress the importance of rigorous evaluation and validation across diverse scenarios before deploying WSI diagnostic models, as an unreliable model could result in catastrophic misdiagnoses in clinical settings.
Appendix A Additional Quantitative Experiments
A.1 Ablation Study
As shown in Table 5, the baseline model mi-Net is very unstable. In our experiments, we found that mi-Net may achieve high training slide-level AUC but extremely low patch-level AUCPR during training and fails completely during testing. This severe overfitting issue disappeared when the dropout was set above 0.4. However, this “training failure” reappeared when training mi-Net using mini-batch gradient descent, since establishing correct classification boundaries for salient and hard instances within a batch of multiple slides is more difficult. FocusMIL does not incorporate dropout; instead, learning the distribution helps to avoid “training failure” and achieves a slide AUC that is 1.3% higher than mi-Net with a dropout of 0.4. Further significant improvements of 7.0% in slide AUC are obtained by training multiple bags together in a mini-batch. Since our method does not rely on bag-inherited information $\bm{Z}_{B}$ for classification, a higher slide AUC actually indicates better classification boundaries for hard instances, which is also confirmed by visualization results.
Table 5: Ablation study on Camelyon16 Dataset. Dropout is set at 0.4 for the second model.
@@ A.1 Ablation Study
FocusMIL 0.871 $\pm$ 0.011 0.400 $\pm$ 0.011
A.2 Ablation Study on Hyperparameter
Ablation Study on KL Divergence Coefficient
As shown in Figure 5, we still observe training failure when $\beta=0$. However, when $\beta$ is set to 0.001, 0.01, or 0.1, the model achieves the desired performance. This further emphasizes the importance of regularizing the latent space. When $\beta=1$, the performance starts to decline, indicating that excessive regularization negatively impacts classification performance.
Figure 5: Results of ablation experiments on KL divergence coefficient and batch size. We report the mean of five experiments
Ablation Study on Batch Size
Batch size greater than 1 can significantly improve the model’s slide-level classification performance. A batch size that is too large can reduce the performance of the model, and we recommend choosing a batch size between 2 and 7.
A.3 Out-of-Distribution Generalization
To further verify that the max-pooling based MIL methods can effectively learn causal factors $\bm{Z}_{C}$, we conduct out-of-distribution (OOD) generalization experiments. Our experimental setup is consistent with CausalMIL (Zhang et al., 2022b), so for the results of the comparison methods, we directly refer to it’s paper.
As shown in Table 6, the three max-pooling-based MIL methods achieve desirable results in both the training and test sets, indicating that they successfully learn the causal factor.
Table 6: Performance on distributionally biased ColoredMNIST dataset.
Method Train Test
ERM $0.849\pm 0.002$ $0.105\pm 0.007$
ERM1 $0.848\pm 0.002$ $0.109\pm 0.005$
ERM2 $0.850\pm 0.002$ $0.101\pm 0.002$
ROBUST MIN MAX $0.843\pm 0.004$ $0.152\pm 0.025$
IRM $0.593\pm 0.044$ $0.628\pm 0.096$
IRM GAME $0.634\pm 0.011$ $0.599\pm 0.027$
iCaRL $0.706\pm 0.008$ $0.688\pm 0.007$
MIVAE $0.810\pm 0.005$ $0.156\pm 0.003$
mi-Net $0.930\pm 0.005$ $0.914\pm 0.008$
CausalMIL $0.919\pm 0.002$ $0.892\pm 0.002$
FocusMIL $0.956\pm 0.010$ $0.935\pm 0.007$
Model Para. Time C16 AUC C17 AUC
ABMIL 657k 3.35s 0.8052 0.7877
DSMIL 856k 4.92s 0.7733 0.7069
TransMIL 2.66M 9.56s 0.8352 0.6899
DTFD-MIL 987k 6.02s 0.8619 0.8192
CausalMIL 302k 3.28s 0.8092 0.5104
FocusMIL 167k 3.23s 0.8706 0.8719
@@ A.3 Out-of-Distribution Generalization
Appendix B Computational Cost Analysis
As shown in Table 7, FocusMIL has significantly fewer parameters compared to attention-based methods. Due to the use of mini-batch gradient descent with a batch size of 3, both CausalMIL and FocusMIL achieve superior training speeds. However, CausalMIL requires additional computations for reconstructing loss and the KL divergence loss of the conditional prior distribution, which results in slightly longer training times compared to FocusMIL.
Appendix C Visualization
@@
For fairness of comparison, the models for visualization are not deliberately selected, and both FocusMIL with a train- ing batch size of 3 and DTFD-MIL models achieved a test slide AUC of approximately 0.87. As shown in Figure 6, our model accurately identifies both large and small tumour regions without any false positive predictions. FocusMIL with a training batch size of 1 misses some small tumour regions. For DTFD-MIL, it misses the tumour region in the first slide and some tumour patches in the second slide, and there are many false positive predictions.
Notably, in the fourth slide of Figure 6 and the normal slide in Figure 7, DTFD-MIL assigns high scores to many patches unrelated to the tumour. After zooming in on these patches, we observe that the tissues in these regions are significantly blurrier, which is associated with the slide preparation process. This indicates that DTFD-MIL relies on bag-inherited information $\bm{Z}_{B}$ for predictions. FocusMIL does not have any false-positive predictions for these normal slides.
Figure 8 demonstrates the visualization results of FocusMIL and DTFD-MIL for five tumour slides. FocusMIL with a training batch size of 3 perfectly predicts all tumour regions. When the training batch size was changed to 1, FocusMIL misses some small tumour regions. This result confirms that training with mini-batch gradient descent does help to establish better classification boundaries for salient and hard instances. For DTFD-MIL, its predictions for both large and small tumour areas are suboptimal due to the fact that it can overfit on salient instances.
@@
Appendix D Experimental setup
D.1 Dataset and Pre-processing
Camelyon16 dataset is widely used for metastasis detection in breast cancer (Bejnordi et al., 2017a). The dataset consists of 270 training and 129 testing WSIs, which yield roughly 2.7 million patches at 10× magnification (the second level in multi-resolution pyramid). Pixel-level labels are available for tumour slides. Every WSI is cropped into $256\times 256$ patches without overlap, and background patches are discarded. A patch is labeled positive if it contains $25\%$ or more cancer areas. The numbers of tumorous versus normal patches are imbalanced as the positive patch ratios of positive slide in the training and testing sets of Camelyon16 are approximately 8.8% and 12.7%, respectively. The complete code for pre-processing will be made publicly available.
Camelyon17 dataset consists of 1,000 WSIs from five hospitals, categorized into different slide labels such as Normal, Isolated Tumour Cells, Micro-metastases, and Macrometastases (Bejnordi et al., 2017b). Since this study focuses on a binary classification task, we relabeled the slides into two categories: Normal and Tumour. Due to the absence of labels in the test set, we only used the 500 slides from the training set. Following the setup of AEM (Zhang et al., 2024), to evaluate the model’s OOD generalization performance, we designated 200 slides from the fourth and fifth hospitals as the test set. The remaining 300 slides were split into training and validation sets with an 8:2 ratio.
D.2 Implementation details
The encoder of FocusMIL consists of a neural network with one hidden layer, with ReLU as the activation function. We use AdamW optimizer with an initial learning rate of $1\times 10^{-4}$ to update the model weights during the training. The dimension of the latent factor is set to 35. The mini-batch size for training FocusMIL model is 3. All experiments are conducted 5 times, and the mean and 0.95 confidence intervals (CI) are reported.
For max-pooling-based methods, we directly use the classifier output as the patch-level prediction. For ABMIL, following the DTFD paper (Zhang et al., 2022a), we use the normalized attention scores for patch-level prediction. For DTFD-MIL, we use the instance probability derivation proposed in their paper for patch-level prediction.
For the attention-based methods, we build other models based on their officially released codes and conduct grid searches for key hyperparameters. For ABMIL, we use gated attention with dropout rate of 0.25 to achieve optimal performance. For CausalMIL (Zhang et al., 2022b), both its encoder and decoder are set up as neural networks with hidden layer neurons 128 and two other fully-connected layers are used to carve factorized prior distribution conditioned on the bag information. All the experiments are conducted with a Nvidia RTX4090 GPU. For more details, please refer to our code in the supplementary.
Appendix E Future Work
In this paper, we have demonstrated that max-pooling-based methods, which adhere to the standard multi-instance assumption, can yield more robust results for WSI classification task. Here are several directions for future exploration:
  1. 1.
    Beyond WSI classification. There may be other tasks that inherently conform to the standard multi-instance assumption but still widely adopt MIL methods that violate this assumption. Re-visiting max-pooling-based approaches in those domains could contribute to more robust decision-making.
  2. 2.
    Alleviating overfitting. Exploring alternative strategies to mitigate overfitting in max-pooling-based methods is a promising direction. Additionally, investigating more complex priors beyond multivariate Gaussian distributions is also worth exploring as a potential avenue for improvement.
  3. 3.
    Improving detection of low-malignancy tumour regions. Identifying small, low-malignancy cancerous areas remains a crucial and challenging problem. While this paper introduces mini-batch optimization, other approaches are also worth exploring.